A 1.00kg brick is sliding along on a rough horizontal surface at 15.0m/s .If the brick stops in 4.60s , how much mechanical energy is lost?What happens to this energy?
Answer #1

E = KE2 - KE1 = 0.5*1*15*15 = -112.5 J


lost in the form of heat
Answer #2

apply mechanical energy lost = kE- mgh

= (0.5 *1* 15*15)   - (1*9.8 * (0.5 * 9.8 * 4.6*4.6))

= 903.6 J